# Linear cost mutual information estimation and independence test of similar performance as HSIC

### v1

@@ Abstract
Evaluation of statistical dependencies between two data samples is a basic problem of data science/machine learning, and HSIC (Hilbert-Schmidt Information Criterion) [1] is considered the state-of-art method. However, for size $n$ data sample it requires multiplication of $n\times n$ matrices, what currently needs $\sim O(n^{2.37})$ computational complexity [2], making it impractical for large data samples. We discuss HCR (Hierarchical Correlation Reconstruction) as its linear cost practical alternative of even higher dependence sensitivity in tests, and additionally providing actual joint distribution model by description of dependencies through features being mixed moments, starting with correlation and homoscedasticity, also allowing to approximate mutual information as just sum of squares of such nontrivial mixed moments between two data samples. Such single dependence describing feature is calculated in $O(n)$ linear time. Their number to test varies with dimension $d$ - requiring $O(d^{2})$ for pairwise dependencies, $O(d^{3})$ if wanting to also consider more subtle triplewise, and so on.
@@ I Introduction
Statistical dependencies start with correlation coefficients, defining dependence between expected values of two random variable: their first moments. There are also higher moments, like variance, skewness, kurtosis - which also contribute to statistical dependencies, like homoscedasticity evaluating co-occurrence of high variance e.g. in financial time series.
Figure 1: Proposed HCR independence test and comparison with state-of-art HSIC having much higher complexity, impractical for large samples. We generate shown $n=1000$ size random dataset from Gaussian mixtures independently for both coordinates, and slightly rotate it to introduce dependence. In HCR we first separately normalize coordinates to nearly uniform distribution in $[0,1]$ with CDF/EDF, then model joint density as a chosen degree polynomial, here $m=4$. Using orthonormal basis of Legendre polynomials, coefficients are approximately: expected value, variance, skewness, kurtosis, and their dependencies. Sum of squares of shown $m^{2}$ nontrivial coefficients approximate mutual information, and allows for independence test: by first normalizing multiplying by $\sqrt{n}$, and then testing if they agree with $N(0,1)$ set of variables, where the presented test of extreme coefficients has turned out the most sensitive to dependencies. There is shown probability of exceeding these values with min/max from $m^{2}=16$ independent $N(0,1)$ variables, also both by taking their product using shown Mathematica code, properly indicating dependence with much higher certainty than HSIC, additionally providing its description with mixed moments and model of joint density.
Hierarchical Correlation Reconstruction (HCR, [3, 4]) automatically decomposes dependencies into such mixed moments of two or more variables. Specifically, like in copula theory [5], we start with normalization of all variables to nearly uniform distribution in $[0,1]$ ($\sim$ quantiles) by $\hat{x}=\textrm{CDF}(x)$ or EDF (cumulative/empirical distribution function) separately for each coordinate. Then we model joint density of such normalized variables as a linear combination in product basis:
@@
for bold $\textbf{x}=(x_{1},\ldots,x_{d})$ denoting dimension $d$ vectors, and $B$ chosen basis of mixed moments. For any continuous joint distribution we can approximate it with polynomials as close as we want, the $a_{\textbf{j}}$ coefficients become moments. Using orthonormal basis of polynomials $\int_{0}^{1}f_{k}(x)f_{l}(x)dx=\delta_{kl}$, coefficients can be MSE estimated as $a_{\textbf{j}}=\frac{1}{n}\sum_{\textbf{x}}f_{\textbf{j}}(\textbf{x})$ mean over normalized sample, in $O(n)$ time for $|X|=n$ data sample.
@@
Alternative way to test independence is verifying if we can factorize joint distribution: $P_{X,Y}=P_{X}P_{Y}$?. It is used e.g. in HSIC independence test [1], considered as the state-of-art. However, it requires matrix multiplication for $n\times n$ matrices like $K_{ij}=\exp(-\|\textbf{x}^{i}-\textbf{x}^{j}\|_{2}/2\sigma^{2})$, and the lowest complexity known algorithm for matrix multiplication has $\sim O(n^{2.37})$ complexity [2], making this test impractical for large data samples.
Therefore, we discuss replacement of $\sim O(n^{2.37})$ HSIC independence test, with practical also for large samples: $O(n|B|)$ cost test based on HCR, for $B$ being a chosen set of features as mixed moments to test, e.g. growing $|B|\sim O(d^{2})$ with square of dimension if including only pairwise dependencies, or $O(d^{2})$ if adding triple-wise and so on.
@@
This is initial version of article, with main purpose to introduce the method. We plan to extend benchmarks comparing with HSIC library [^1] HSIC library used for benchmarks: http://pypi.org/project/PyRKHSstats/ and details in the next version.
Figure 2: Comparison of joint density estimation using global basis of polynomials/moments we use in HCR and local basis of Gaussians usually used in HSIC, requiring to choose radius $\epsilon$. Global basis allows to represent the density with a reasonable number of features (moments) e.g. $m^{2}$ here, allowing the discussed complexity reduction. Evaluating log-likelihood in cross-validation, local basis barely got above 0 obtained also for trivial $\rho=1$ assumption, while global basis was able to find features as moments, which generalize well from training to test set. Here with the highest log-likelihood for $m=4$ degree up to kurtosis, we focus on in this article. Intuitively, global features like moments are better at generalization, than local features just assuming that new points will be close to the old points.
@@ II HCR independence test and evaluation
Assume we have $Z=\{(\textbf{x}^{i},\textbf{y}^{i})\}_{i=1}^{n}$ as size $n$ data samples as vectors of $d_{x}$, $d_{y}$ dimensions, for which we would like to test independence. To simplify presentation we focus on $d_{x}=d_{y}=d$ equal dimension case, but also discuss general situation.
In HCR it is convenient to start with normalization of marginal distributions to nearly uniform in $[0,1]$, then model their joint distribution as a linear combination in some orthonormal basis, usually polynomials - which should use zero coefficients for independent variables, we will build independence test from.
@@ II-A Normalization to nearly uniform in $[0,1]$ with CDF/EDF
As in copula theory citecopula, we start with normalization of variables to nearly uniform distribution in $[0,1]$ by transforming with CDF/EDF (cumulative or empirical distribution function) to nearly quantiles, independently for each coordinate.
@@
Figure 3: Top left: the first $j=0,1,2,3,4$ from orthonormal polynomial basis $f_{j}(x)$ for $d=1$ dimension and product basis $f_{j}(x)f_{k}(y)$ for $d=2$ dimensions, which linear combination is used as model of (joint) density in HCR, e.g. $\rho(x,y)=\sum_{jk=1}^{m}a_{jk}f_{j}(x)f_{k}(y)$ for variables normalized to nearly uniform distribution in $[0,1]$. As $f_{0}=1$, top row $a_{i0}$ describes marginal distribution of $x$ and $a_{0j}$ left column of $y$. Then $i,j\geq 1$ describe their dependencies as mixed moments close to expected value, variance, skewness and kurtosis. Right: $d=3$ case adding triplewise dependencies.
Bottom left: standard view on (joint/conditional) entropy of $X,Y$ variables and their mutual information. Right: for HCR using $\ln(1+w)\approx w$ first Taylor term they can be approximated with sums of squares of coefficients, especially mutual information as $I(X;Y)\approx\sum_{i,j=1}^{m}\left(a_{ij}\right)^{2}$, also in higher dimension, which should be close to zero for independent variables.
@@
As mutual information is invariant under reparametrization of the marginal variables [9], we can estimate it for such transformed variables instead: $I(U;V)\approx I(X;Y)$.
@@ II-B Polynomial density model
Alternatively there could be used Fourier or DCT basis, especially for periodic variables.
The coefficients can be interpreted as mixed moments, as $f_{0}=1$ only between nonzero indexes, e.g. $a_{00i0}$ is $i$-th moment of 3rd out of 4 variables, $a_{012}$ describe dependence between expected value of 2nd variable, and variance of 3rd variable.
@@ II-C Entropy, mutual information estimation
Using $\ln(1+t)\approx t$ approximation and orthonormality, we can approximate entropy in nits ($1/\ln(2)\approx 1.44$ bits) from the coefficients:
@@
II-C1 Subtracting the variance
Assume we get $n$ values $\{w^{i}\}_{i=1..n}$ from $N(\mu,\sigma)$ distribution, and we want to estimate $\mu^{2}$. A natural approach is calculating the mean $\bar{w}=\frac{1}{n}\sum_{i}w_{i}$ and taking $\bar{w}^{2}$, however, it would be artificially increased.
@@ II-C1 Subtracting the variance
Applying square to both sides, and taking expected values: mean square from $N(0,1)$ is 1, hence we get:
$n\,E[(\bar{w}-\mu)^{2}/\sigma^{2}]=1$
@@
$I_{c}(\hat{X},\hat{Y})=\sum_{\mathbf{j}\in B_{x}^{+}}\ \sum_{\mathbf{k}\in B_{y}^{+}}\left(\textrm{mean}(f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y}))^{2}-\frac{1}{n}\textrm{var}(f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y}))\right)$ (11)
@@ II-D HCR-based independence tests
To test it, let us first normalize them as in CLT:
@@
Assuming $H_{0}$ hypothesis, for pairwise dependencies: between one coordinate of both samples, due to independence and normalization above ’var’ variance has to be 1, allowing to omit the above estimation from sample. However, for higher order dependencies it does not have to be true due to dependencies inside samples - requiring to estimate variance from sample.
Accordingly to CLT, for $H_{0}$ independence hypothesis, all $\hat{a}_{\textbf{j},\textbf{k}}$ should be from approximately $N(0,1)$.
@@
II-D1 Log-likelihood test
of $\textrm{mean}(\ln(\textrm{PDF}(\hat{a}_{\textbf{j},\textbf{k}})))$, which assuming $H_{0}$ and large $M$ is approximately from normal distribution. For $N(0,1)$ PDF we get $E[\ln(\textrm{PDF})]=-(1+\ln(2))/2\approx-1.4189$ and variance 1/2, leading to test:
@@ II-D1 Log-likelihood test
Figure 4: $\textrm{PDF}^{M}_{i}$ probability densities of sorted $M=1,\ldots,16$ variables from $N(0,1)$ normalized Gaussian distribution, which can be used for tests if estimated normalized coefficients are from $N(0,1)$ for $\mathcal{H}_{0}$ hypothesis of independence, especially for the extreme values indicating the main dependence direction.
II-D2 Testing extremes $\{\hat{a}_{\textbf{j},\textbf{k}}\}_{\textbf{j}\in B_{x}^{+},\textbf{k}\in B_{y}^{+}}$
@@
For hypothesis testing, we need to find distribution for $M$ sorted values from $N(0,1)$. To find $\textrm{PDF}_{M,i}(x)$ density of $i$-th in order, there are i1i-1 smaller or equal, and M1iM-1-i greater then equal, combinatorially leading to $M{M-1\choose i-1}$ possibilities, with density, shown in Fig. 4:
@@
using incomplete Euler Beta function. We can use $\textrm{CDF}_{M,i}$ to test probability of obtaining one or a few extreme values.
II-E Choice of basis $B$

### v2

@@ Abstract
Evaluation of statistical dependencies between two data samples is a basic problem of data science/machine learning, and HSIC (Hilbert-Schmidt Information Criterion) [1] is considered the state-of-art method. However, for size $n$ data sample it requires multiplication of $n\times n$ matrices, what currently needs $\sim O(n^{2.37})$ computational complexity [2], making it impractical for large data samples. We discuss HCR (Hierarchical Correlation Reconstruction) as its linear cost practical alternative, in tests of even higher sensitivity to dependencies, and additionally providing actual joint distribution model for chosen significance level, by description of dependencies through features being mixed moments, starting with correlation and homoscedasticity. Also allowing to approximate mutual information as just sum of squares of such nontrivial mixed moments between two data samples. Such single dependence describing feature is calculated in $O(n)$ linear time. Their number to test varies with dimension $d$ - requiring $O(d^{2})$ for pairwise dependencies, $O(d^{3})$ if wanting to also consider more subtle triplewise, and so on.
@@ I Introduction
Statistical dependencies start with correlation coefficients, defining dependence between expected values of two random variable: their first moments. There are also higher moments, like variance, skewness, kurtosis - which also contribute to statistical dependencies, like homoscedasticity evaluating co-occurrence of high variance common e.g. in financial time series or Fig. 1.
Figure 1: Proposed HCR independence test and comparison with state-of-art HSIC of much higher complexity, impractical for large samples, which are required to find very subtle dependencies. We generate shown $n=1000$ size random dataset from Gaussian mixtures independently for both coordinates, and slightly rotate it to introduce dependence. In HCR we first separately normalize coordinates to nearly uniform distribution in $[0,1]$ by CDF/EDF, then model joint density as a chosen degree polynomial, here $m=4$. Using orthonormal basis of Legendre polynomials, coefficients are approximately: expected value, variance, skewness, kurtosis, and their dependencies. Sum of squares of shown $m^{2}$ nontrivial coefficients approximate mutual information, and allows independence test: by first normalizing multiplying by $\sqrt{n}$, and then testing if they agree with $N(0,1)$ set of variables, where the presented test of extreme coefficients has turned out the most sensitive to dependencies, also evaluating significance of contributions to joint distribution. There is shown probability of exceeding these values with min/max from $m^{2}=16$ independent $N(0,1)$ variables, also both by taking their product using shown Mathematica code, properly indicating dependence with much higher certainty than HSIC, additionally providing its description with mixed moments and model of joint density.
Hierarchical Correlation Reconstruction (HCR, [3, 4]) automatically decomposes dependencies into such mixed moments of two or more variables. Specifically, like in copula theory [5], we start with normalization of all variables to nearly uniform distribution in $[0,1]$ ($\sim$ quantiles) by $\hat{x}=\textrm{CDF}(x)$ or EDF (cumulative/empirical distribution function), separately for each coordinate. Then we model joint density of such normalized variables as a linear combination in product basis:
@@
for bold $\textbf{x}=(x_{1},\ldots,x_{d})$ denoting dimension $d$ vectors, and $B$ chosen basis of mixed moments. For any continuous joint distribution we can approximate it with polynomials as close as we want, the $a_{\textbf{j}}$ coefficients become moments. Using orthonormal basis of (Legendre) polynomials $\int_{0}^{1}f_{k}(x)f_{l}(x)dx=\delta_{kl}$, coefficients can be MSE estimated as $a_{\textbf{j}}=\frac{1}{n}\sum_{\textbf{x}}f_{\textbf{j}}(\textbf{x})$ mean over normalized sample, in $O(n)$ time for $|X|=n$ data sample.
@@
Alternative way to test independence is verifying if we can factorize joint distribution: $P_{X,Y}=P_{X}P_{Y}$? It is used e.g. in HSIC independence test [1], considered as the state-of-art. However, it requires matrix multiplication for $n\times n$ matrices like $K_{ij}=\exp(-\|\textbf{x}^{i}-\textbf{x}^{j}\|_{2}/2\sigma^{2})$, and the lowest complexity known algorithm for matrix multiplication has $\sim O(n^{2.37})$ complexity [2], making this test impractical for large data samples.
Therefore, we discuss replacement of $\sim O(n^{2.37})$ HSIC independence test, with practical also for large samples: $O(n|B|)$ cost test based on HCR, for $B$ being a chosen set of features as mixed moments to test, e.g. growing $|B|\sim O(d^{2})$ with square of dimension if including only pairwise dependencies, or $O(d^{3})$ if adding very subtle triplewise and so on.
Beside much lower cost, making it practical to search very large data samples required to find subtle dependencies, HCR also turns out more sensitive to dependencies. It additionally provides approximation of mutual information, and model of joint distribution with control of significance of contributions.
@@
This is early version of article, with main purpose to introduce the method. We plan to extend benchmarks comparing with HSIC library [^1] HSIC library used for benchmarks: http://pypi.org/project/PyRKHSstats/ and details in future versions.
Figure 2: Comparison of joint density estimation using global basis of polynomials/moments we use in HCR, and local basis of Gaussians usually used in HSIC, requiring to choose radius $\epsilon$. Global basis allows to represent the density with a reasonable number of features (moments) e.g. $m^{2}$ here, allowing the discussed reduction of complexity e.g. from $\sim O(n^{2.37})$ to $O(m^{2}n)$ for $n$ point independence test. Evaluating log-likelihood in cross-validation, local basis barely got above 0 obtained for trivial $\rho=1$ assumption, while global basis finds features as moments, which generalize well from training to test set. Here with the highest log-likelihood for $m=4$ degree up to kurtosis, quite universal e.g. for financial data, we focus on in this article. In contrast, local basis just assumes new points will be close to the old ones, what does not generalize well.
Figure 3: Top left: the first $j=0,1,2,3,4$ from orthonormal polynomial basis $f_{j}(x)$ for $d=1$ dimension and product basis $f_{j}(x)f_{k}(y)$ for $d=2$ dimensions, which linear combination is used as model of (joint) density in HCR, e.g. $\rho(x,y)=\sum_{jk=1}^{m}a_{jk}f_{j}(x)f_{k}(y)$ for variables normalized to nearly uniform distribution in $[0,1]$. As $f_{0}=1$, top row $a_{i0}$ describes marginal distribution of $X$ and $a_{0j}$ left column of $Y$. Then $i,j\geq 1$ describe their dependencies as mixed moments close to expected value, variance, skewness and kurtosis. Right: $d=3$ case adding much more subtle triplewise dependencies.
Bottom left: standard view on (joint/conditional) entropy of $X,Y$ variables and their mutual information. Right: for HCR using $\ln(1+t)\approx t$ first Taylor term they can be approximated with sums of squares of coefficients, especially mutual information as $I(X;Y)\approx\sum_{i,j=1}^{m}\left(a_{ij}\right)^{2}$, also in higher dimensions, which should be close to zero for independent variables.
@@ II HCR independence test and evaluation
Assume we have $X=\{\textbf{x}^{i}\}_{i=1}^{n},Y=\{\textbf{y}^{i}\}_{i=1}^{n}$ size $n$ data sample as pairs forming $Z=\{(\textbf{x}^{i},\textbf{y}^{i})\}_{i=1}^{n}$ of vectors of $d_{x}$, $d_{y}$ dimensions, for which we would like to test independence. To simplify presentation we focus on $d_{x}=d_{y}=d$ equal dimension case, but also discuss general situation.
In HCR it is convenient to start with normalization of marginal distributions to nearly uniform in $[0,1]$, then model joint distribution as a linear combination in some orthonormal basis, usually polynomials - which for independent samples should use zero nontrivial moments between them, we can build test from.
@@ II-A Normalization to nearly uniform in $[0,1]$ with CDF/EDF
As in copula theory  [5], we start with normalization of variables to nearly uniform distribution in $[0,1]$ by transforming with CDF/EDF (cumulative or empirical distribution function) to nearly quantiles, independently for each coordinate.
@@
[…]
@@ II-B Polynomial density model
The coefficients can be interpreted as mixed moments, only between nonzero indexes as $f_{0}=1$, e.g. $a_{00i0}$ is $i$-th moment of 3rd out of 4 variables, $a_{012}$ describes dependence between expected value of 2nd variable, and variance of 3rd variable.
Alternatively there could be used Fourier or DCT basis, especially for periodic variables. As mentioned further, the basis can be automatically optimized e.g. with SVD [10], for given data sample, or family of data.
@@ II-C Entropy, mutual information estimation
Using $\ln(1+t)\approx t$ first order approximation and orthonormality, we can approximate entropy in nits ($1/\ln(2)\approx 1.44$ bits) from the coefficients:
@@
Subtracting the variance
Assume we get $n$ values $\{w^{i}\}_{i=1..n}$ from $N(\mu,\sigma)$ distribution, and we want to estimate $\mu^{2}$. A natural approach is calculating the mean $\bar{w}=\frac{1}{n}\sum_{i}w_{i}$ and taking $\bar{w}^{2}$, however, it would be biased (artificially increased).
@@
Applying square to both sides, and taking expected values: mean square of value from $N(0,1)$ is 1, hence we get:
$\frac{n}{\sigma^{2}}\,E[(\bar{w}-\mu)^{2}]=1$
@@
$I_{c}(\hat{X},\hat{Y})=\sum_{\mathbf{j}\in B_{x}^{+}}\ \sum_{\mathbf{k}\in B_{y}^{+}}\left(\textrm{mean}(f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y}))^{2}-\frac{\textrm{var}(f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y}))}{n}\right)$ (11)
@@
Also, while we have used only the first term of Taylor series $\ln(1+t)=-\sum_{i=1}^{\infty}(-t)^{i}/i$ , it is worth to somehow include higher terms, requiring integrals of 3 and more basis functions.
@@ II-D HCR-based independence tests
To test it, let us first normalize them as in CLT, so that for $H_{0}$ independence hypothesis, all $\hat{a}_{\textbf{j},\textbf{k}}$ should be from $N(0,1)$:
@@
Assuming $H_{0}$ hypothesis, for pairwise dependencies: between one coordinate of each samples, due to independence and normalization above ’var’ variance has to be 1, allowing to omit the above estimation from sample. However, for higher order dependencies it does not have to be true due to dependencies inside samples - requiring to estimate variance from sample.
@@
Log-likelihood test
of $\textrm{mean}(\ln(\textrm{PDF}(\hat{a}_{\textbf{j},\textbf{k}})))$, which assuming $H_{0}$ and large $M$ is approximately from $N(0,1)$, for which we get $E[\ln(\textrm{PDF})]=-(1+\ln(2))/2\approx-1.4189$ and variance 1/2, leading to test:
@@
Figure 4: $\textrm{PDF}^{M}_{i}$ probability densities of sorted $M=1,\ldots,16$ variables from $N(0,1)$ normalized Gaussian distribution (analogously for different distributions), which can be used for tests if estimated normalized coefficients are from $N(0,1)$ for $\mathcal{H}_{0}$ hypothesis of independence, especially for the extreme values indicating the main dependence direction. As in Fig. 5, it also allows to estimate significance of individual contributions, e.g. for joint distribution model of chosen significance level.
Testing extremes $\{\hat{a}_{\textbf{j},\textbf{k}}\}_{\textbf{j}\in B_{x}^{+},\textbf{k}\in B_{y}^{+}}$
@@
For hypothesis testing, we need to find distribution of sorted $M$ values from independent $N(0,1)$ or a different single variable distribution having given $\textrm{PDF}=\textrm{CDF}^{\prime}$, e.g. for Marchenko-Pastur for SVD optimization. To find $\textrm{PDF}_{M,i}(x)$ density of $i$-th in order, there are i1i-1 smaller or equal, and M1iM-1-i greater then equal, combinatorially leading to $M{M-1\choose i-1}$ possibilities, with density, shown in Fig. 4:
@@
using incomplete Euler Beta function. We can use $\textrm{CDF}^{M}_{i}$ to test sorted values for hypothesis of being from a given distribution, also find those disagreeing for a given significance level.
II-E Chosen statistical significance joint distribution model
[…]
$\min\left(\textrm{CDF}^{M}_{i}(\textrm{sort}(\hat{a})_{i}),\ 1-\textrm{CDF}^{M}_{i}(\textrm{sort}(\hat{a})_{i})\right)<\alpha$ (16)
for $N(0,1)$ distribution in joint density formula (3), or Marchenko-Pastur-like for SVD optimized as in Fig. 5.
Figure 5: Top: we can perform SVD (singular value decomposition) of $|B_{x}^{+}|\times|B_{y}^{+}|$ matrix of coefficients, allowing to optimize their bases (18), leading to decomposition of modeled joint density into contributions of controlled significance, allowing to use the highest or a few below chosen significance level $\alpha$, here leading to similar Pr(independent) as HCRboth in Fig. 1. To distinguish signal from noise, we could use random matrix theory for singular values of $|B_{x}^{+}|\times|B_{y}^{+}|$ matrix with $N(0,1)$ random coefficients. Marchenko-Pastur theorem [13] handles such situations, however, for the limit of infinite size, e.g. bounding singular values by 4. In contrast, the plots shows densities from $10^{6}$ generated $4\times 4$ random matrices of $N(0,1)$ coefficients, clearly exceeding this bound - we can use such empirical estimation for practical approximation. It allows for joint distribution models using statistically significant contributions, e.g. for 0.01 significance we should take normalization plus the first two. Bottom: analogously directly using coefficients for various significance level using (16) formula. There are also shown $\textrm{PDF}^{M}_{i}$ densities for $M=16$ sorted $N(0,1)$ values, and sorted $\hat{a}$ coefficients being $N(0,1)$ for independence hypothesis, showing that non-extreme values can also be significant.
II-F Choice of basis $B$